# https://www.vrvis.at

### v1

@@ 4.2 Continual Learning
[…]
@@
Table 3: Results for CL experiments. Performance is measured on all classes after the final CL step (i.e., after 4 steps). Our method is on-par with CL baselines while learning substantially faster.
ID Train data (#samples) Method & Setup $F$ pre-train/train $\tau_{p}$ [h] $\tau_{l}$ [s] $\text{mIoU}_{\text{cl}}$ $\text{mIoU}_{\text{cat}}$
B0 CITY, fine (2975) JOINT (upper bound) no / yes 0 20.62 81.7 86.6
B1 CITY, fine (2975) FT (lower bound) no / yes 0 12.76 4.0 4.3
B2 CITY, fine (2975) REPLAY, $m\!=\!256$ no / yes 0 5.11 56.2 64.2
B3 CITY, fine (256) GDUMB, $m\!=\!256$ no / yes 0 110.9 74.4 80.3
B4 CITY, fine (256) ours B0 ($1^{\text{st}}$ step) / no 3.07 0.32 66.0 70.7
B5 CITY, fine (256) ours B3 ($1^{\text{st}}$ step) / no 1.42 0.32 65.4 71.0
B6 CITY, fine (256) GDUMB, $m\!=\!256$ A9 / yes 2.41 63.13 76.9 82.1
B7 CITY, fine (256) ours A9 / no 2.41 0.32 75.4 81.3
[…]
REPLAY (B2) GDUMB (B3) ours (B4) ours (B5) GDUMB (B6) ours (B7)
[…]
4.3 Few-Shot Learning
For FSL, we pre-train $F$ on coarse labels (A9). We then perform experiments with 10, 8, 6, 4, and 2 fine-labeled support samples. FSL via FT is performed by fine-tuning $\theta$ on support samples. To study FT under realistic conditions, we divide the support sets into 2 splits with the latter containing 4 samples used for early stopping. For FSL with PARMESAN, we fill $M$ with all support samples. A low number of support samples can lead to a high variance in performance. We thus repeat each experiment 5 times with randomly drawn support samples, where we fix the random generation process across experiments to ensure comparability. We report average performances on the full validation set in Tab. 5.
Table 5: Results for FSL experiments, averaged over 5 repetitions. We use a feature extractor pre-trained on coarse labels and a support set containing fine labels. Our method learns fast and is data-efficient especially for small support sets.
ID Train data (#samples) Method & Setup $F$ pre-train/train $\tau_{p}$ [h] $\tau_{l}$ [s] $\text{mIoU}_{\text{cl}}$ $\text{mIoU}_{\text{cat}}$
C0 CITY, fine (10, train on 6) FT A9 / yes 2.41 32.29 71.7 76.4
C1 CITY, fine (8, train on 4) FT A9 / yes 2.41 69.53 71.1 75.9
C2 CITY, fine (6, train on 2) FT A9 / yes 2.41 225.6 69.5 74.7
C3 CITY, fine (10) ours, $m\!=\!10$, $\phi\!=\!1$ A9 / no 2.41 0.17 72.2 78.2
C4 CITY, fine (8) ours, $m\!=\!8$, $\phi\!=\!1$ A9 / no 2.41 0.21 71.1 77.5
C5 CITY, fine (6) ours, $m\!=\!6$, $\phi\!=\!1$ A9 / no 2.41 0.26 69.9 76.2
C6 CITY, fine (4) ours, $m\!=\!4$, $\phi\!=\!1$ A9 / no 2.41 0.34 66.9 73.3
C7 CITY, fine (2) ours, $m\!=\!2$, $\phi\!=\!1$ A9 / no 2.41 0.61 58.3 65.6
FSL with PARMESAN can be done up to 370 times faster w.r.t. FT (C2, C7) and remains on par in terms of predictive performance for equal support set sizes (C0-C5). Our method is data-efficient in the sense that is able to utilize the entire support set at inference and that it can be applied in combination with a small support set (C6, C7).
4.4 Limitations
Depending on the choice of hyperparameters such as $m$ and $\phi$, PARMESAN might require more computational resources and longer inference times compared to learned task-heads. Memory consumption and inference speed can potentially be further improved by introducing sparsity to stored feature pyramids, i.e., reducing the number of unique patterns per memory sample. Moreover, choosing a suitable feature extractor might require taking a trade-off between a general-purpose model versus a data- and task-specific model.
@@ 5 Conclusion
In this work we proposed PARMESAN, a flexible, parameter-free approach that separates computation from memory and leverages transduction for solving dense prediction tasks. We also proposed a message passing approach to take advantage of intra-query correlations. Our method can perform complex tasks such as CL and FSL in an easy, intuitive, and optionally private manner without any continuous training of learnable parameters.
We demonstrated that our approach learns substantially faster than common baselines while achieving competitive predictive performance, showing stable knowledge retention, and being data-efficient.
Future Work: Possible future work includes leveraging self-supervised feature extractors, studying OOD detection capabilities, as well as investigating memory-related topics such as dataset distillation. Performance might be further increased by implementing algorithmic or methodical improvements. This includes, e.g., using explorative memory search instead of our current greedy approach, as well as applying novelty search instead of random sampling for memory consolidation. Finally, PARMESAN might be beneficial to other fields such as reinforcement learning, where fast adaptation can be important.
@@
Philip Matthias Winter, Maria Wimmer, David Major, Dimitrios Lenis, Astrid Berg, Theresa Neubauer, Gaia Romana De Paolis, Johannes Novotny, Sophia Ulonska, Katja Bühler
@@ Appendix 0.A Method
Initialization: We now describe the initialization of $G$ in more detail. First, we need to initialize hyperparameters such as the number of data dimensions $dim$, the number of memory samples $m$ and the resolutions $(res)^{l}$ for all levels. We then define connectivity kernels between nodes of different levels, which remain fixed for all samples. Each node in level $l$ has a parent node in level $l+1$. For the highest level $l=n$, all nodes have a single root node as their parent. Each node in level $l$ typically has several child nodes in level $l-1$, e.g., $n_{ch}=2^{dim}$. Leaf nodes in level $l=0$ refer to memory labels. We assume that resolutions of $h^{1}$, $x$, and $z$ match, implying that each node in level $l=1$ has exactly one child node in $l=0$. We initialize parent indices $i_{k,par}$ such that hierarchical search has access to the full memory content. We initialize parent similarity scores $s_{k,par}$ with a value of 1.0 for all nodes.
@@
We implemented all methods and experiments using the PyTorch framework [63]. We apply the exact same pre- and postprocessing to retain comparability. We conduct all experiments on a single NVIDIA GeForce RTX 3090 (24 GB) graphics card. PARMESAN has inference times in the order of 1 second per image for $m=256$, $\phi=0.5$, and an input resolution of 512$\times$1024 pixels. By default, we store memory content on RAM and transfer representations to GPU when needed for hierarchical search. In case of small GPU RAM, the hierarchical search can be chunked into smaller pieces via windowing and performed sequentially without compromising performance. For MP, all operations can be done on GPU.
Appendix 0.B Experiments
@@ 0.B.1 Datasets
From the Cityscapes semantic segmentation benchmark dataset (CITY) [7], we use the 5000 images with fine semantic and disparity annotations, and the 19998 images with coarse semantic annotations (see Fig. 5). We use the official splits for training and validation with 2975 and 500 samples, respectively. Since labels for the 1525 test samples are not publicly available, we report best performances achieved on the validation split. For JSRT [84], a 80:10:10 split is applied, resulting in 199, 24, and 24 samples for training, validation, and testing, respectively. For CITY, we use 34 semantic classes with 8 categories for training and the official 19 classes for evaluation. For JSRT, we use 6 classes for training and evaluation: background, left lung, right lung, heart, left clavicle, and right clavicle. We rescale all samples to a resolution of 512$\times$1024 pixels for CITY and 512$\times$512 pixels for JSRT.
For CITY disparity maps, we observe that some pixels near image borders are labeled incorrectly. We manually mask out these erroneous labels and treat them as invalid measurements.
@@ 0.B.2 Models, Methods, Training and Evaluation
Models: We use a U-Net [77] variant with residual blocks [30] and instance normalization [96] as our default architecture, resulting in 9 levels (excluding the label-level $l=0$) with respective feature map resolutions of [512[512[ 512, $256$, $128$, $64$, $32$, $16$, $8$, $4$, 2]2]2 ] pixels per dimension. The respective number of feature maps are [6[6[ 6, $12$, $24$, $48$, $96$, $192$, $384$, $768$, 1536]1536]1536 ], resulting in a total of 73.15m learnable parameters.
For ablations studies, we additionally use a ConvNeXt Tiny with pre-trained weights for feature extraction [50]. Pre-trained weights were obtained by training ADE20K [112] for semantic segmentation. The model comprises 4 levels with feature map resolutions of $[128,64,32,16]$ and corresponding number of feature maps $[96,192,384,786]$. To be able to use it with our transduction module, we complete the feature pyramid, i.e., add missing resolutions, via interpolation and extrapolation of neighboring representations. We reduce the number of feature maps in high resolutions by a factor of 4 to meet our hardware limitations (see paragraph “Training“ below) and keep the number of feature maps constant for missing lower resolutions. This results in $[6,24,96,192,384,768,768,768,768]$ feature maps at the respective resolutions of $[512,256,128,64,32,16,8,4,2]$.
Method: For memory search, we set $n_{ch}=2^{dim}$ for all levels. For MP, we set $\lambda=1$, $k=16$. Convergence is typically achieved within 16 MP steps. For TTA, we apply three separate forward passes and down-scale inputs by factors $s=\{0.8,0.9,1\}$, where $s=1$ implies no transformation.
[…]
Evaluation: We use exponential moving average models [65, 79, 89] with a rate of 0.99 for evaluation. Early stopping [58] with a patience of 3 and a validation verbosity of 3000 update steps is used (400 for JSRT). Early stopping in CL experiments only depends on the performance on available classes at the respective CL step. We use a validation verbosity of 30 update steps for FSL FT experiments due to small support sets.
@@
Results: We show qualitative prediction results for all ablation studies, CL experiments, and FSL experiments in Fig. 6, Fig. 7, and Fig. 8, respectively.
@@ 0.B.3 Continual Learning Setup
[…]
In contrast to image classification, samples can contain classes that belong to different CL steps, which is referred to as overlapping labels. For the CI-CL scenario, this requires introducing a dedicated background class to avoid data-leakage between different CL steps. Therefore, non-available classes are re-labeled as background in each CL step. This implies that the background class is different for different CL steps. As proposed in recent work [5], we ignore background pixels for validation.
@@
In our experiments, we apply classical replay (REPLAY) as a baseline method. In order to keep labels consistent for replay training, we employ an adaptive masking approach. For samples of the current CL step $i$, we mask out all background pixels, i.e., pixels which do not belong to currently active classes. In contrast, for memory samples we mask out all pixels which belong to currently active classes, i.e., keeping pixels of old classes and the background. Training minibatches for REPLAY consist of standard samples in addition to replay samples, which we select randomly from $M$. For REPLAY, we observe that using one or more replay samples from $M$ per update step leads to strong memory overfitting, which in turn leads to reduced validation performance. We achieve best results with one replay sample every other minibatch.
@@
Figure 8: Model predictions for all FSL experiments with two different query inputs.

### v2

@@ 4.2 Continual Learning
We perform CI-CL experiments, where the goal is to incrementally learn new semantic classes while at the same time automatically inferring the given task. We employ two different CL scenarios: 2-2 (4) cat and 13-1 (7) cl [95]. At 2-2 (4) cat, data at every CL step comprises labels for all classes of different categories: $\textit{D}_{1}$: void & flat, $\textit{D}_{2}$: construction & object, $\textit{D}_{3}$: nature & sky, and $\textit{D}_{4}$: human & vehicle. For memory-based methods REPLAY, GDUMB, and PARMESAN, we update samples in $M$ with labels from the current CL step while keeping the stored inputs. REPLAY and GDUMB utilize $M$ for training $\theta$, whereas PARMESAN uses $M$ for transduction. We summarize our results in Table 2, Table 3, Fig. 3, and provide additional information in Section 0.B.3.
FT (B1, B12) and JOINT (B0, B11) behave as expected with strong forgetting and superior performance w.r.t. other methods, respectively. For same memory sizes, PARMESAN is on par with other CL baselines while enhancing flexibility, allowing simple learning and unlearning of individual examples, and boosting learning speed between 3-4 orders of magnitude. We further demonstrate that our method has stable knowledge retention properties (see Table 3). This effect can be attributed to our shift towards memory-induced forgetting, where we explicitly state what to learn (new classes), what to retain (old classes), and what to forget (background).
@@
Table 2: Results for CL experiments on CITY fine, including size of $M$ and learning times $\tau_{l}$. Performance is measured on all classes after the final CL step.
ID scenario Method & Setup $M$ [GB] $\theta$ pre-train $\theta$ train $\tau_{l}$ [s] $\text{mIoU}_{\text{cl}}$ $\text{mIoU}_{\text{cat}}$
B0 2-2 (4) cat JOINT - A6 3.64 81.6 86.6
B1 2-2 (4) cat FT - A6 3.44 4.1 4.3
B2 2-2 (4) cat REPLAY, $m\!=\!940$ 6.43 A6 2.79 76.8 82.6
B3 2-2 (4) cat GDUMB, $m\!=\!940$ 6.43 A6 10.7 78.9 84.5
B4 2-2 (4) cat MiB - A6 6.10 78.8 83.9
B5 2-2 (4) cat ours, $m\!=\!256$ 0.55 A6 0.012 73.9 79.9
B6 2-2 (4) cat ours, $m\!=\!940$ 2.02 A6 0.010 75.6 81.3
B7 2-2 (4) cat ours 6.40 A6 0.012 76.9 82.4
B8 2-2 (4) cat ours, $m\!=\!256$, ConvNeXt 0.55 [45] 0.021 71.6 78.3
B9 2-2 (4) cat ours, $m\!=\!940$, ConvNeXt 2.0 [45] 0.019 73.3 80.0
B10 2-2 (4) cat ours, ConvNeXt 6.35 [45] 0.021 78.2 84.3
B11 13-1 (7) cl JOINT - A6 3.97 80.4 85.2
B12 13-1 (7) cl FT - A6 1.89 0.21 0.21
B13 13-1 (7) cl REPLAY, $m\!=\!940$ 6.43 A6 7.72 79.1 84.1
B14 13-1 (7) cl GDUMB, $m\!=\!940$ 6.43 A6 9.36 79.1 84.1
B15 13-1 (7) cl MiB - A6 2.44 65.2 72.5
B16 13-1 (7) cl ours 6.40 A6 0.001 76.8 82.2
B17 13-1 (7) cl ours, ConvNeXt 6.35 [45] 0.003 78.1 84.0
Table 3: Knowledge retention after 4 CL steps, showing absolute changes $\delta$ between initial and final $\text{mIoU}_{\text{cl}}$. Datasets $\textit{D}_{i}$ comprise semantic classes of different categories according to the respective CL step. Our method has stable knowledge retention and can not suffer from parameter-induced forgetting by design.
REPLAY (B2) GDUMB (B3) MiB (B4) ours (B6) ours (B7) ours (B9) ours (B10)
$\textit{D}_{i}$ initial $\delta$ initial $\delta$ initial $\delta$ initial $\delta$ initial $\delta$ initial $\delta$ initial $\delta$
$\textit{D}_{1}$ 78.70 -2.80 77.89 -0.71 79.73 -2.60 71.21 -0.79 72.05 -1.03 64.39 -0.95 67.99 -0.45
$\textit{D}_{2}$ 49.66 +3.62 55.87 +0.07 57.21 -3.29 49.05 -0.43 50.52 -0.51 50.17 -0.50 54.40 -0.43
$\textit{D}_{3}$ 63.05 +4.24 69.78 -0.38 69.12 -1.65 61.68 -0.25 63.12 -0.26 59.27 -0.30 62.47 -0.26
$\textit{D}_{4}$ 39.60 - 48.37 - 46.67 - 40.33 - 42.03 - 49.27 - 54.68 -
$avg$ 57.75 +1.69 62.98 -0.25 63.18 -2.51 55.57 -0.49 56.93 -0.60 55.78 -0.58 59.88 -0.38
$std$ 16.95 3.90 13.32 0.39 14.35 0.82 13.62 0.27 13.29 0.39 7.31 0.33 6.57 0.10
@@ 5 Conclusion
Summary: We propose PARMESAN, a flexible, parameter-free transduction approach for dense prediction tasks. We also propose a message passing approach to take advantage of intra-query correlations, as well as a novelty-sparsity approach to retain memory efficiency and diversity. Our method is capable of performing CL in an easy and intuitive manner without any continuous training of learnable parameters. Our approach was demonstrated to learn substantially faster than common baselines while achieving competitive predictive performance and showing stable knowledge retention properties.
Limitations: Depending on hyperparameters like $m$, $\phi$, and $n_{sp}$, PARMESAN can require more computational resources than learned task-heads. Moreover, choosing a suitable feature extractor requires taking a trade-off between a general-purpose model versus a data- and task-specific model.
Open Problems: Possible research opportunities include leveraging self-supervised feature extractors, studying explainability aspects, as well as investigating memory-related topics such as dataset distillation for memory consolidation. Inference speed and memory requirements might be further improved, e.g., by using specialized libraries. Finally, PARMESAN can be beneficial to other fields such as domain adaptation, transfer learning, and few-shot learning, where flexibility and fast learning are often important.
@@
Philip Matthias Winter*, Maria Wimmer*, David Major, Dimitrios Lenis, Astrid Berg, Theresa Neubauer, Gaia Romana De Paolis, Johannes Novotny, Sophia Ulonska, Katja Bühler
@@ Appendix 0.A Method
Initialization: We now describe the initialization of $G$ in more detail. First, we need to initialize hyperparameters such as the number of data dimensions $dim$, the number of memory samples $m$ and the resolutions $(res)^{l}$ for all levels. We then define connectivity kernels between nodes of different levels, which remain fixed for all samples. Each node in level $l$ has a parent node in level $l+1$. For the highest level $l=n$, all nodes have a single root node as their parent. Each node in level $l$ typically has several child nodes in level $l-1$, e.g., $n_{ch}=2^{dim}$. Leaf nodes in level $l=0$ refer to memory labels $z_{j}$. We assume that resolutions of $h^{1}$, $x$, and $z$ match, implying that each node in level $l=1$ has exactly one child node in $l=0$. We initialize parent indices $i_{k,par}$ such that hierarchical search has access to the full memory content. We initialize parent similarity scores $s_{k,par}$ with a value of 1.0 for all nodes.
@@
We implemented all methods and experiments using the PyTorch framework [58]. We apply the exact same pre and postprocessing to retain comparability. We conduct all experiments on a single NVIDIA GeForce RTX 3090 (24 GB) graphics card. PARMESAN has inference times in the order of 1 second per image for $m=256$, $\phi=0.5$, and an input resolution of 512$\times$1024 pixels. By default, we store memory content on RAM and transfer representations to GPU when needed for hierarchical search. In case of small GPU RAM, the hierarchical search can be chunked into smaller pieces via windowing and performed sequentially without compromising performance. For MP, all operations can be done on GPU.
Appendix 0.B Experimental Setup and Additional Experiments
In this section, we provide more details regarding our overall experimtal setup and perform some additional experiments.
@@ 0.B.1 Datasets
From the Cityscapes semantic segmentation benchmark dataset (CITY) [8], we use the 5000 images with fine semantic and disparity annotations, and the 19998 images with coarse semantic annotations (see Fig. 5). We use the official splits for training and validation with 2975 and 500 samples, respectively. Since labels for the 1525 test samples are not publicly available, we report best performances achieved on the validation split. For JSRT [76], a 80:10:10 split is applied, resulting in 199, 24, and 24 samples for training, validation, and testing, respectively. For CITY, we use 34 semantic classes with 8 categories for training and the official 19 classes for evaluation. For JSRT, we use 6 classes for training and evaluation: background, left lung, right lung, heart, left clavicle, and right clavicle. We rescale all samples to a resolution of 512$\times$1024 pixels for CITY and 512$\times$512 pixels for JSRT. For CITY disparity maps, we observe that some pixels near image borders are labeled incorrectly. We manually mask out these erroneous labels and treat them as invalid measurements.
@@ 0.B.2 Models, Methods, Training and Evaluation
Models: We use a U-Net [69] variant with residual blocks [26] and instance normalization [87] as our default architecture, resulting in 9 levels (excluding the label-level $l=0$) with respective feature map resolutions of $[512\times 1024,256\times 512,128\times 256,64\times 128,32\times 64,16\times 32,8\times 16,4\times 8,2\times 4]$ pixels. The respective number of feature maps are [6[6[ 6, $12$, $24$, $48$, $96$, $192$, $384$, $768$, 1536]1536]1536 ], resulting in a total of 73.15m learnable parameters.
For some experiments, we use a ConvNeXt Tiny with pre-trained weights for feature extraction [45]. Pre-training was performed on ADE20K [99] for semantic segmentation. We use the publicly available, pre-trained weights provided by [45]. The ConvNeXt model comprises 4 levels with feature map resolutions of $[128,64,32,16]$ and corresponding number of feature maps $[96,192,384,786]$. To be able to use it with our transduction module, we complete feature pyramids, i.e., add missing resolutions, via interpolation and extrapolation of neighboring representations. For ConvNeXt, we use the same number of feature maps at respective resolutions as for U-Net.
Method: For memory search, we set $n_{ch}=4^{dim}$ for all levels. For MP, we set $\lambda=1$, $k=16$. Convergence is typically achieved within 32 MP steps. For test-time augmentation (TTA), we use 3 forward passes with down-scaled and reflection-padded variants of $x_{q}$. Registration of down-scaled variants with the original variant is achieved by up-scaling and center-cropping the respective predictions. We down-scale inputs by factors $s=\{0.8,0.9,1\}$, where $s=1$ implies no transformation. A final prediction is obtained by averaging over all passes.
Training: Some of our experiments (baseline methods and pre-training for PARMESAN) require training of learnable parameters $\theta$. We use the cross-entropy loss as the default loss function for semantic segmentation. For monocular depth estimation, we regress $d:=\textit{log}\textit{(depth)}$) and use the scale-invariant loss [13] with $\lambda=0.5$. We mask out invalid labels both during training and evaluation. Training is performed via stochastic gradient descent [66], utilizing the backpropagation algorithm [33, 70] and the AdamW optimizer [34, 47] with default hyperparameters, a minibatch size of 8, a constant learning rate of 0.0001, and a weight decay of 0.00001. We apply gradient-norm clipping [57], allowing maximum gradient norms of 50. We train our models on randomly cropped patches of 512$\times$512 pixels for CITY and on full images of 512$\times$512 pixels for JSRT. For data augmentation during training, we use horizontal flips ($p=0.5$, not for JSRT), vertical flips ($p=0.5$, not for JSRT), rotations ($p=0.5$, ±$\pm 25^{\circ}$), Gaussian noise ($p=0.5$, $\sigma=0.05$), as well as changes in brightness, saturation, and contrast.
Evaluation: We use exponential moving average models [59, 71, 82] with a rate of 0.99 for evaluation. Early stopping [53] with a patience of 5 and a validation verbosity of 3000 update steps is used (400 for JSRT). Early stopping in CL experiments only depends on the performance on available classes at the respective CL step. Since classes are highly imbalanced in the 13-1 (7) cl scenario, we apply early stopping using both old and current classes for JOINT and GDUMB.
@@
Results: We show qualitative prediction results for all ablation studies and CL experiments experiments in Fig. 6 and Fig. 7, respectively. Experiments A14, A15, and A16 are not mentioned in the main paper text. Results in Table 4 indicate moderate memory requirements for our method. Performance remains robust in case of low $m$ (A14) or when using less channels (A15, A16). Iterative MP and sequentially executed TTA lead to higher $\tau_{i}$ w.r.t. our best setup (A1 / A2 / A3).
Table 4: Size of $M$ (samples stored on GPU), peak allocated GPU memory $\mathit{C}$ at inference, inference speed $\tau_{i}$, and predictive performance on CITY.
ID Method & Setup $M$ [GB] $\mathit{C}$ [GB] $\tau_{i}$ [s] $\tau_{l}$ [s] $\text{mIoU}_{\text{cl}}$ $\text{mIoU}_{\text{cat}}$
A0 JOINT - 0.93 0.0068 3.72 82.5 87.4
A1 ours 6.40 11.0 1.43 0.012 82.1 87.0
A2 ours, no MP 6.40 10.9 0.25 0.012 81.4 86.4
A3 ours, no TTA 6.40 10.8 0.48 0.012 81.2 86.3
A14 ours, $m\!=\!4,\phi\!=\!1$ 0.009 4.44 1.25 0.011 72.8 78.1
A15 ours, 75% channels 4.82 9.10 1.37 0.011 82.0 86.9
A16 ours, 50% channels 3.29 7.40 1.32 0.010 81.8 86.7
@@ 0.B.3 Continual Learning Setup
The goal in CL is learning a sequence of $N$ tasks $\mathcal{T}=((T_{1},\textit{D}_{1}),...,(T_{N},\textit{D}_{N}))$ on respective datasets $\mathcal{D}=(\textit{D}_{1},...,\textit{D}_{N})$. Old data $\textit{D}_{j}$ is inaccessible for CL step $i$ for $j<i$. For memory-based CL methods, data from $\textit{D}_{j}$ can still be retained if it is stored in $M$. Given a task-solving model (”solver”) $S:X\rightarrow Z$ with inputs $x\in X$, predictions $y\in Z$, and labels $z\in Z$, the objective is minimizing the unbiased sum of losses $\mathbb{E}_{x,z}[L(y,z)]_{\mathcal{T}}$ over all tasks on the entire data distribution.
We employ two different CL scenarios: 2-2 (4) cat and 13-1 (7) cl. At 2-2 (4) cat, data at every CL step comprises labels for all classes of different categories: $\textit{D}_{1}$: void & flat, $\textit{D}_{2}$: construction & object, $\textit{D}_{3}$: nature & sky, and $\textit{D}_{4}$: human & vehicle. As proposed by Yang et al. [95], the 13-1 (7) cl scenario starts with 13 classes at CL step $i=1$ (i.e., road, sidewalk, building, wall, fence, pole, traffic light, traffic sign, vegetation, terrain, sky, person, rider) and adds one new class at every following CL step (i.e., car, truck, bus, train, motorcycle, bicycle).
In contrast to image classification, samples for semantic segmentation can contain classes that belong to different CL steps, which is referred to as overlapping labels. For the CI-CL scenario, this requires introducing a dedicated background class to avoid data-leakage between different CL steps. Therefore, non-available classes are re-labeled as background in each CL step. This implies that the background class is different for different CL steps. As proposed in recent work [5], we ignore background pixels for validation.
@@
In order to keep labels consistent for REPLAY, we employ an adaptive masking approach. For samples of the current CL step $i$, we mask out all background pixels, i.e., pixels which do not belong to currently active classes. In contrast, for memory samples we mask out all pixels which belong to currently active classes, i.e., keeping pixels of old classes and the background. Training minibatches for REPLAY consist of standard samples in addition to replay samples, which we select randomly from $M$. For REPLAY, we observe that using 4 or more replay samples from $M$ per update step leads to strong memory overfitting, which in turn leads to reduced validation performance. For $m=940$, we achieve best results with minibatches composed of 3 replay samples and 8 standard samples.
For MiB and REPLAY baselines, the background class is learned in the first CL step. During successive CL steps, background is partially unlearned again depending on new classes.